# High-mobility heavy quasiparticles in a van der Waals antiferromagnetic dense Kondo lattice CeTe3

### v1

@@ ACKNOWLEDGMENTS
This research was funded by the National Key R&D Program of China (2023YFA1609600 and 2022YFA1602602), the National Natural Science Foundation of China (U23A20580, 12174180, 12274155), the Guangdong Basic and Applied Basic Research Foundation (2022B1515120020), and the Beijing National Laboratory for Condensed Matter Physics (2024BNLCMPKF004). We thank the HM (https://cstr.cn/31125.02.SHMFF.HM) at the Steady High Magnetic Field Facility, CAS (https://cstr.cn/31125.02.SHMFF), for providing technical support and assistance in data collection and analysis.
@@
Y.L. conceived and designed the experiments. H.Z. grew the crystal and performed most of the bulk-property measurements with the aids from S.Z., Z.W., K.L. and Y.Y. C.D. and J.W. carried out the dHvA measurements with pulsed-field magnetization facility. J.C., J.Z., C.X. measured the SdH quantum oscillations under static high magnetic field. B.J. and Y.D. made the optical conductivity measurements. Y.Z. and J.L. fabricated the nanoflakes and performed the SdH measurements on the devices. K.W. provided constructive suggestions. H.Z., Y.D., J.L. and Y.L. discussed the data, interpreted the results, and wrote the paper with input from all the authors.
@@
High-mobility heavy quasiparticles in a van der Waals antiferromagnetic dense Kondo lattice
Hai Zeng1∗, Yang Zhang1∗, Bingke Ji2∗, Jiaqiang Cai3, Shuo Zou1, Zhuo Wang1, Chao Dong1, Kangjian Luo1, Yang Yuan1, Kai Wang4, Jinglei Zhang3, Chuanyin Xi3, Junfeng Wang1, Yaomin Dai2†, Jing Li1‡, and Yongkang Luo
1 Wuhan National High Magnetic Field Center and School of Physics, Huazhong University of Science and Technology, Wuhan 430074, China
2 School of Physics, Nanjing University, Nanjing 210023, China
3 Anhui Key Laboratory of Low-Energy Quantum Materials and Devices, High Magnetic Field Laboratory, Hefei Institutes of Physical Science, Chinese Academy of Sciences, Hefei 230031, China
@@ 2 School of Physics, Nanjing University, Nanjing 210023, China
School of Physics and Electronics, Henan University, Kaifeng 475004, China
@@
In this Supplemental Information (SI), we provide detailed information about methods, crystalline characterization, magnetization, magnetoresistance, optical spectroscopy, Hall effect and two-band fitting, and LK formula fitting of quantum oscillations.
SI \@slowromancapi@: XRD, magnetization, and magnetoresistance
@@
(A) X-ray diffraction pattern of a CeTe3 single crystal, where only $(0,2k,0)$ diffractions are observed. The lattice parameters $a$, $b$, and $c$ at room temperature are 4.24(9), 23.25(4), and 4.52(3) Å, respectively. The large ratio $b/a(c)\sim 5.5$ indicates a quasi-two-dimensional character of crystal structure.
(B) Temperature-dependent magnetic susceptibility of bulk CeTe3 for different crystallography directions. The navy solid line represents the Curie-Weiss fit of the inverse powder-averaged magnetic susceptibility, $1/\chi_{avg}=3/(\chi_{b}+2\chi_{ac})$. The fitting yields effective magnetic moment $\mu_{eff}$ = 2.62 $\mu_{B}$ and the Weiss temperature $\theta_{W}=-35.11$ K.
@@
SI \@slowromancapii@: Optical spectroscopy data of CeTe3
@@
(B) Real part of the optical conductivity, $\sigma_{1}(\omega)$. Signatures for multiple CDW gaps are observed near 5000 cm-1 for all the measured temperatures and 2800 cm-1 for below 200 K. Inset shows $\sigma_{1}(\omega)$ at representative 300 and 5 K in the spectral range 50 - 50000 cm-1.
Figure S3. Temperature dependent $1/\tau_{1}$ from optical conductivity. On the whole, $1/\tau_{1}$ is very small, characteristic of low scattering rate and high charge mobility. An upturn in $1/\tau_{1}$ is observed near $T_{\text{K}}^{\text{on}}$ due to the onset of incoherent Kondo scattering. The drop of $1/\tau_{1}$ near 200 K is likely owing to the CDW order.
SI \@slowromancapiii@: Magnetotransport of CeTe3
Figure S4. Magnetotransport of CeTe3.
(A) $\mathbf{H}\parallel\mathbf{b}$ MR of bulk CeTe3. The inset shows the MR curve at 1.5 K up to 34 T.
@@
(C) Hall conductivity $\sigma_{xy}$ of bulk sample and its fitting to a two-band model.
(D) ibid, but for nanoflaked D3 sample.
@@
(A) $n_{e,h}$ of bulk CeTe3.
(B) $\mu_{e,h}$ of bulk CeTe3.
(C) $n_{e,h}$ of CeTe3 nanoflacke D3.
(D) $\mu_{e,h}$ of CeTe3 nanoflacke D3.
SI \@slowromancapiv@: Additional quantum oscillation data

### v2

@@ ACKNOWLEDGMENTS
This research was funded by the National Key R&D Program of China (2023YFA1609600 and 2022YFA1602602), the National Natural Science Foundation of China (U23A20580, 12174180 and 12274155), the Guangdong Basic and Applied Basic Research Foundation (2022B1515120020), and the Beijing National Laboratory for Condensed Matter Physics (2024BNLCMPKF004). We thank the HM (https://cstr.cn/31125.02.SHMFF.HM) at the Steady High Magnetic Field Facility, CAS (https://cstr.cn/31125.02.SHMFF), for providing technical support and assistance in data collection and analysis.
@@
Y.L. conceived and designed the experiments. H.Z. grew the crystals and performed most of the bulk-property measurements with the aids from S.Z., Z.W., K.L. and Y.Y. C.D. and J.W. carried out the dHvA measurements with pulsed-field magnetization facility with the technical support of L.L. J.C., J.Z., C.X. measured the SdH quantum oscillations under static high magnetic field. B.J. and Y.D. made the optical conductivity measurements. Y.Z. and J.L. fabricated the nanoflakes and performed the SdH measurements on the devices. K.W. provided constructive suggestions. H.Z., Y.D., J.L. and Y.L. discussed the data, interpreted the results, and wrote the paper with input from all the authors.
@@
Kondo-coupled van der Waals antiferromagnet with high-mobility quasiparticles
Hai Zeng1∗, Yang Zhang1∗, Bingke Ji2∗, Jiaqiang Cai3, Shuo Zou1, Zhuo Wang1, Chao Dong1, Kangjian Luo1, Yang Yuan1, Kai Wang4, Jinglei Zhang3, Chuanyin Xi3, Junfeng Wang1, Liang Li1,5, Yaomin Dai2†, Jing Li1‡, and Yongkang Luo1,5,6§
1 Wuhan National High Magnetic Field Center and School of Physics, Huazhong University of Science and Technology, Wuhan 430074, China;
2 School of Physics, Nanjing University, Nanjing 210023, China;
3 Anhui Key Laboratory of Low-Energy Quantum Materials and Devices, High Magnetic Field Laboratory, Hefei Institutes of Physical Science, Chinese Academy of Sciences, Hefei 230031, China;
@@ 2 School of Physics, Nanjing University, Nanjing 210023, China;
School of Physics and Electronics, Henan University, Kaifeng 475004, China;
@@
State Key Laboratory of Advanced Electromagnetic Technology, Huazhong University of Science and Technology, Wuhan 430074, China;
6
@@
In this Supplemental Information (SI), we provide additional results that will further support the discussions and conclusion in the main text, including crystalline characterization, magnetization, magnetoresistance, optical spectroscopy, Hall effect and two-band fitting, and Lifshitz-Kosevich analysis of quantum oscillations.
SI I: XRD, magnetization, and magnetoresistance
@@
(A) X-ray diffraction pattern of a CeTe3 single crystal, where only $(0,2k,0)$ diffractions are observed. The lattice parameters $a$, $b$ and $c$ at room temperature are 4.24(9), 23.25(4) and 4.52(3) Å, respectively. The large ratio $b/a(c)\sim 5.5$ indicates a quasi-two-dimensional character of the crystal structure.
(B) Temperature-dependent magnetic susceptibility of bulk CeTe3 for different crystallography directions. The navy solid line represents the Curie-Weiss fit of the inverse powder-averaged magnetic susceptibility, $1/\chi_{\text{avg}}=3/(\chi_{b}+2\chi_{ac})$. The fitting yields effective magnetic moment $\mu_{\text{eff}}$ = 2.62 $\mu_{\text{B}}$ and the Weiss temperature $\theta_{\text{W}}=-35.11$ K.
@@
SI II: Optical spectroscopy data of CeTe3 and LaTe3
@@
(B) Real part of the optical conductivity, $\sigma_{1}(\omega)$. Signatures for multiple CDW gaps are observed near 5000 cm-1 for all the measured temperatures and 2800 cm-1 for below 200 K. Inset shows $\sigma_{1}(\omega)$ at representative 300 and 5 K in the full spectral range 50 - 50000 cm-1.
Figure S3. Temperature dependent $1/\tau_{1}$ extracted from optical conductivity.
On the whole, $1/\tau_{1}$ is very small, characteristic of low scattering rate and high charge mobility. An upturn in $1/\tau_{1}$ is observed near $T_{\text{K}}^{\text{on}}$ due to the onset of incoherent Kondo scattering. The drop of $1/\tau_{1}$ near 200 K is likely owing to the CDW order.
Figure S4. Optical spectroscopy data of LaTe3.
(A) Real part of the optical conductivity, $\sigma_{1}(\omega)$. Signatures for multiple CDW gaps are observed near 6000 cm-1 and 3000 cm-1.
(B) Left, temperature dependence of $\omega_{\text{p}}^{2}\equiv\omega_{\text{p1}}^{2}+\omega_{\text{p2}}^{2}$; right, temperature dependent average effective mass of charge carriers, $\langle m^{*}\rangle$. The inset shows the sum of hole and electron carrier concentrations. Note that upon cooling, $\omega_{\text{p}}^{2}$ increases monotonically at low temperature, which stands in stark contrast to that of CeTe3. This leads to the great difference in $\langle m^{*}\rangle$ between LaTe3 and CeTe3.
SI III: Magnetotransport of CeTe3
Figure S5. Magnetotransport data of CeTe3.
(A) $\mathbf{H}\parallel\mathbf{b}$ magnetoresistance (MR) of bulk CeTe3. The inset shows the MR curve at 1.5 K up to 34 T.
@@
(C) Hall conductivity $\sigma_{xy}$ of bulk sample, and fitting to a two-band model.
(D) Hall conductivity $\sigma_{xy}$ of nanoflake sample D3, and fitting to a two-band model.
@@
(A) $n_{\text{e,h}}$ of bulk CeTe3.
(B) $\mu_{\text{e,h}}$ of bulk CeTe3.
(C) $n_{\text{e,h}}$ of CeTe3 nanoflake D3.
(D) $\mu_{\text{e,h}}$ of CeTe3 nanoflake D3.
SI IV: Additional quantum-oscillation data of CeTe3
@@
Figure S8. Additional evidence for the topological transition across $T_{\text{N}1}$ in CeTe3.
(A-D) Landau fan index diagram of sample S2 at selected temperatures.
(E) Temperature dependent intercept of the Landau fan plot. The hatched area indicates the AFM1 transition near $T_{\text{N}1}$. A jump of Berry phase $\Phi_{\text{B}}$ from $\pi$ to 0 is clearly seen near $T_{\text{N}1}$. Sample S2 was measured under field up to 16 T, while sample S1 up to 34 T (shown in Figure 3).